OpenStudy (anonymous):

"An object is projected vertically upward from ground level with velocity v. Find the height in meters attained by the object." Can anybody solve this, presumably in terms of Galileo's formula? I know the answer is the initial velocity squared divided by 2g (due to it being at the back of the book), but I have no idea how we get there.

OpenStudy (anonymous):

decelration due to gravity v-gt=0 gives t=v/g 2nd formula s=ut+at^2/2 =v*v/g-g*(v/g)^2/2=v^2/2g

OpenStudy (anonymous):

After some staring and head-scratching, I got it. Thanks a lot, vvijay!